OpenStudy (anonymous):

Write an equation of a line with a slope of 0 and passing through the point (5,4).

OpenStudy (amtran_bus):

Point slope formula \[y-y1=m(x-x1)\] That is, \[y-4=0(x-5)\] so \[y-4=0\] adding 4 to both sides gives y=4

OpenStudy (anonymous):

thank you

OpenStudy (amtran_bus):

Welcome!